# Synthesizing data products, mathematical models, and observations for lake temperature forecasting

### v2

@@ 3.4 Modeling forecast bias
Now, let $\hat{\mu}_{\_}n^{\mathrm{SK}}(\mathbf{x})$ and $\hat{\sigma}^{\mathrm{SK}}(\mathbf{x})^{2}$ denote the moments of that NOAA-GLM surrogate (Eq. 11), combining Sections 3.23.3, but fit to the entire campaign of runs across all 31 ensemble members and all training days/horizons/depths under study: ($\mathbf{X}_{\_}{n_{\_}M},\mathbf{Y}_{\_}{n_{\_}M}$), and $n_{\_}M\approx$ 8 million-odd runs. Providing $\mathbf{x}\in\mathcal{X}$ representing any time point(s) of interest (future, present or past), including depths and horizons of interest, yields a prediction (or forecast) capturing uncertainty in GLM dynamics driven by NOAA ensemble hindcasts. Recall from Section 2.2 that estimating a bias correction requires measuring discrepancy between computer model predictions and field data observations, forming a data set $D^{b}$ that can be used to train a GP. Our field data are a single set of observations $\mathbf{Y}^{F}_{\_}{n_{\_}F}$ in time and depth, coded in $\mathbf{X}_{\_}{n_{\_}F}$. In Section 2.2 we had $\mathbf{X}_{\_}{n_{\_}F}=\mathbf{X}_{\_}{n_{\_}M}$, so $\mathbf{Y}^{F}_{\_}{n_{\_}F}$ could easily be lined up with predictions $\boldsymbol{\mu}_{\_}{n_{\_}M}$ to measure discrepancies (6). However, now our set of GLM inputs (and outputs) is expanded [Section 3.1] to encompass horizon, which is not relevant to our field sensor measurements. Field data and surrogate are out of alignment.
@@
Recall that $\hat{\mu}_{\_}n^{\mathrm{SK}}(\mathcal{X})$ is the predictive mean of the first moments’ fit $\mathcal{GP}(D_{\_}n^{(m)})$. As in Section 2.2, $\sigma_{\_}n^{2}(\mathcal{X})$ over-estimates the variance of the sum of two GP variances because it is missing a negative covariance term. In contrast to the discussion in Section 2.2, we cannot utilize full predictive covariance matrices $\Sigma(\mathcal{X})$ because the Vecchia implementation we use only provides the diagonal, which is what we are denoting as $\sigma_{\_}n^{2}(\mathcal{X})$. We may put in for $\mathcal{X}$ anything we wish; however, the most useful settings indicate hindcasts ($\mathcal{X}$ depicting a historical set of forecasting days, horizons, and depths) and forecasts ($\mathcal{X}$ denoting the current day, a horizon of the next thirty days, and all depths of interest).
@@
To illustrate, Figure 6 shows examples of 1m-depth “slices” of hindcasts via quantities defined above starting on Feb. 19, 2021 (left), 2022 (middle) and 2023 (right). Consider the top row first. Each thin gray line is a NOAA-GLM hindcast over a 30-day horizon. Thicker, orange lines summarize the computer model surrogate ($\hat{\mu}_{\_}n^{\mathrm{SK}}$ and $\hat{\sigma}_{\_}n^{\mathrm{SK}}$ via 90% error bars), denoted as GPGLM. Darker dots show the $\bar{\mathbf{Y}}^{F}_{\_}{n_{\_}F}$ corresponding to that date/depth and each horizon. Discrepancy data $D^{b}$, derived from the $\hat{\mu}_{\_}n^{\mathrm{SK}}$ and $\bar{\mathbf{Y}}^{F}_{\_}{n_{\_}F}$, are shown at the bottom of each panel. Observe that in all three panels/years, these fits exhibit a cold bias because the bias correction is positive. The strength and statistical significance of this conclusion varies considerably across the three years. In 2021 correction from zero might not be statistically significant. But in 2022 and 2023 it is substantial (upwards of 3 degrees), highly non-linear, and error-bars do not cover zero over most of the 30-day horizon.
The bottom row of the figure shows how Eq. (12) is used to combine the quantities in the top and middle row, showing the bias fit, to obtain a bias-corrected surrogate (GPBC). Observe that in all three years this is a good approximation to sensor measurements. Note that while PIs for the surrogate in the top panels exhibit a trumpet-like shape over increasing horizon, those in the bottom panels no longer have this pronounced shape. GPBC is more confident than GPGLM because the latter uses $\mathbb{V}\mathrm{ar}\{Y(\mathbf{\mathbf{x}})\}$ but the former leverages $\hat{\sigma}^{\mathrm{SK}}_{\_}n(\mathbf{x})^{2}$ from Eq. (11). The total variance of GPBC is obtained by adding $\hat{\sigma}^{\mathrm{SK}}_{\_}n(\mathbf{x})^{2}$ to the predictive variance from the bias GP $(\sigma_{\_}n^{2b}(\mathcal{X}))$, shown in Eq. (12). Because $\hat{\sigma}^{\mathrm{SK}}_{\_}n(\mathbf{x})^{2}\ll\sigma_{\_}n^{2b}(\mathcal{X})$ from the (homoskedastic) bias GP, the overall predictive variance of GPBC appears homoscedastic.
@@
Figure 7 shows a similar pair of views, but now in a forecasting context corresponding to the right-most column of Figure 6 (2023). Specifically, we mean that the gray lines, from the GLM were used to build the surrogate (along with simulations for all other days/depths in the past), but the field data (dark dots) are not observed. Thus, bias correction is based only on patterns learned on historical data. Although the overall predictive quality in Figure 7 is lower than the right column of Figure 6, because the former is in-sample and the later is out-of-sample, it is clear the GPBC performs well. Visually, the bias correction is effective (panel B) compared to the surrogate on its own (panel A). This can also be quantified as follows: GPBC achieves an RMSE of 1.24C compared to 1.86C GPGLM on its own. Note that while PI coverage is lower for GPBC for this particular example, the width of the PIs of GPBC (2.23) is less than half of the PI width of the surrogate on its own (5.14).
@@ 4.1 Iterative forecasting
Our forecasting framework consists of: an initial training stage, in which many historical NOAA-GLM simulations are used to fit a GP surrogate $\mathcal{GP}(D_{\_}{n_{\_}{M}})$ and bias correction $\mathcal{GP}(D^{b}_{\_}n)$; and an iterative stage, in which new data are downloaded and a new forecast is generated from updated surrogates. Here we are using $\mathcal{GP}$ generically, since many variations are possible. The initial training phase begins with obtaining NOAA-GLM blue hindcasts for all reference dates in the historical training period. A surrogate fit to these data provides a statistical abstraction of NOAA-GLM simulations over the historical period of time. Observed residuals from sensor data are be used to estimate a bias correction. The iterative phase, which happens daily, is more interesting. This begins with downloading NOAA ensemble forecasts over the next thirty days. Next, these are fed as driver data to the GLM, augmenting the simulation campaign from the training phase, and updating the surrogate and bias correction accordingly. The bias-corrected surrogate can then be used to make lake temperature forecasts over that thirty-day period. The process repeats the next day after observing actual temperatures on the current day from sensors at each depth. Additional detail, along with a diagrammatic depiction [Figure 14] is provided in Appendix B.
@@ 4.2 Incorporating annual variability into forecasts
The inputs to our surrogate modeling and bias correction framework are Julian day $t\in\{1,\dots,365\}$, depth $d\in\{0,\dots,9\}$, and horizon $h\in\{1,\dots,30\}$. A consequence of this choice, especially with regard to the time input $t$, is that our forecasts will be identical for any day $t$ in any year. We have a “day of year” model. As we gather more data, forecasts will have more conservative UQ because they will have been trained on a wider diversity of observations across year for that day, and days nearby in time. In other words, they will better capture climatological variability. But they will also be less reactive to current weather patterns and other environmental variables.
Figure 8: In-sample forecasts for March 1, 2023 models without $\phi$ (“without $\phi$”; left panels, dotted lines) and those with $\phi$ (“with” $\phi$); right panels, solid lines) at 1m depth; 90% PIS are denoted by dashed lines, Sensor measurements as purple dots. GLM forecasts for March 1, 2023 are in dark gray; previous years in light gray.
To illustrate, consider the left column of Figure 8, which shows examples of predictions over a 30-day horizon for March 1, 2023 at depth 1m from the surrogate only (top panel) and after bias correction (bottom). In contrast to previous visuals, we now show all three sets of NOAA-GLM simulations available up until this time point: today’s (March 1 2023) in black and the previous two years (March 1 in 2022 and 2021) in gray. Observe that the forecasts do a poor job of tracking today’s black simulations. In the top row, forecasts from the surrogate track the corpus of simulations, whereas in the bottom row their spread is narrowed by the bias correction, but in an overly conservative fashion.
Toward obtaining more dynamic forecasts, we introduce a fourth input variable, $\phi$, whose behavior is illustrated in the right panels in the figure. With $\phi$ we aim to capture the unique “state” for day $t$ and year $\lambda$ [first introduced in Section 3.1], via the following calculation
@@
i.e., average sensor-measured temperatures near the surface (depths $d$ = 0 and 1 meters) over the previous five days. Temperatures near the surface are more variable than those at deeper layers, and are therefore a good indication of weather conditions. We augment design matrices $\bar{\mathbf{X}}_{\_}n$ and $\bar{\mathbf{X}}_{\_}{n_{\_}F}^{F}$ to include $\phi(t,\lambda)$ matching up with $t$ and $\lambda$. (Recall that the $\lambda$ column is ignored in our GP modeling, but now with the addition of $\phi$ its presence is felt through the state that it captures (Eq. 13)). Surrogate modeling and bias correction, as described in Sections 3.23.4, are not affected by this change excepting higher input dimension and additional lengthscale(s). In particular, responses $\bar{\mathbf{Y}}_{\_}n^{M}$ and $\mathbf{Y}^{F}_{\_}{n_{\_}F}$ remain as before.
The right column of Figure 8 explores the effect of adding $\phi$ as a predictor. Observe how, in contrast with the left column without $\phi$, the GLM surrogate without a bias-correction (top/orange) provides a good fit to NOAA-GLM forecasts (black lines). When bias-correction is added (bottom/orange), prediction intervals narrow considerably compared to the sensor measurements (purple dots). We generally find better results with the addition of $\phi$, as demonstrated in Section 5. Consequently, we shall going forward take the $\phi$ version to be the main/default modeling variation unless we mention that modeling “without $\phi$” was entertained as a contrast. In fact, all previous visuals in the paper used the version with $\phi$, because they looked better. Eq. (13) represents only one of many ways that a year-differentiating state could be calculated. It could even be treated as a latent variable [see Appendix E].
@@ 5.1 Validation and benchmarking
When new sensor measurements come in at midnight we have the opportunity to assess the accuracy of yesterday’s forecasts, as predicted on all thirty previous days at varying future horizons. We evaluate (1) RMSE; (2) a proper log score (Eq. (25) of Gneiting and Raftery, 2007); (3) confidence interval width and (4) coverage. For details see Appendix C. Smaller RMSEs and larger scores are preferred.
@@
    raw 1–30 day-ahead NOAA-GLM forecasts in the testing time period. We calculated empirical means and standard deviations over all 31 ensemble members for all combinations reference date, depth, and horizon in the test time period.
@@ 5.2 Forecasting exercise
Panel (C) summarizes empirical coverage from 90% predictive intervals over horizon. Only the methods using surrogate-assisted NOAA-GLM forecasts can achieve nominal coverage (at 90%) at any horizon. Raw-GLM dramatically under-covers. Its surrogate (GPGLM), by contrast, is able to correctly quantify uncertainty by accessing the entire corpus of simulations across the years. Although empirical coverage drops off from nominal at higher horizons, models that incorporate GLM forecasts (GPBC and GPGLM) always outperform OGP, the climatological benchmark. Additional insight into these coverage results is provided by panel (D). Observe that the PIs of raw GLM forecasts are much narrower than those from other competitors, especially at early horizons. We also know (e.g., Figure 6) that standalone GLM simulations and forecasts can exhibit pronounced bias. Taken together, this explains raw GLM’s poor scores (panel B). Also note that PIs resulting from GPGLM are twice as wide as those of GLM for all horizons. Although GPBC and GPGLM have similar PI coverage, PIs of GPBC are much narrower than those of GPGLM. In fact, at horizon 30, PI width of GPBC (2.2) is nearly half that of GPGLM (4.2).
Figure 10: Forecasts from GPGLM (blue) and GPBC (orange) generated during the forecasting exercise for horizon 1 (one day into the future, top) and horizon (30, 30 days into the future, bottom) over day of year (DOY) at surface, 2, 4, 6, and 8m depths (each panel denotes one depth). Observations are shown in darker purple dots.
Appendix D.1 provides a similar analysis over depth. Figure 10 focuses on forecasts over day of year (DOY) for two forecast horizons (1 and 30 days into the future). The top panel shows one-day ahead forecasts generated by GPGLM and GPBC, while the bottom panel shows forecasts at 30 days into the future at five depths. Observe in the top panels (one-day ahead) that GPGLM and GLM are generally in agreement with observations, lying nearly on top of one another. At thirty days out (bottom panels), the discrepancies between forecasts from GPGLM and GPBC are much more apparent, especially for deeper depths. At surface depths, the effect of bias-correction is not as obvious, but the improvement from bias-correction enhances accuracy at deeper depths between DOY 1 and 240.
Figure 11: Bias (e.g., $\bar{\mathbf{Y}}_{\_}n^{F}-\hat{\mu}_{\_}n^{\mathrm{SK}}(\bar{\mathbf{X}}_{\_}n)$ for out-of sample forecasts from the GLM surrogate (solid pink line), raw GLM (solid gray line) and predicted bias (dot-dash teal line), averaged over depth and horizon for dates in the testing time period. Vertical black dashed lines denote example DOYs depicted in Figure 16.
Next, Figure 11 shows how raw GLM bias (gray) surrogate GLM bias (“actual bias” via $\bar{\mathbf{Y}}_{\_}n^{F}-\hat{\mu}_{\_}n^{\mathrm{SK}}(\bar{\mathbf{X}}_{\_}n)$), and predicted bias vary over DOY in our out-of-sample forecasting exercise. Observe that while accuracy varies over horizon [Figure 9] the magnitude of bias from GLM depends even more strongly on the time of year a forecast is made [Figure 11]. Bias from GLM and its surrogate are similar. Predicted bias, which is used to correct it, follows a similar pattern, but it is not perfectly aligned with actual bias: it underestimates for the first half of the year, and then overestimates in the second half. Note that the predicted bias is not smooth, because bias for each day in the forecasting period was predicted iteratively, so each prediction was based upon a slightly different GP fit. Appendix D.2 provides an additional set of visuals breaking out horizon and depth contributions to bias for three particular days.

### v3

@@ 3.4 Modeling forecast bias
Now, let $\hat{\mu}_{\_}n^{\mathrm{SK}}(\mathbf{x})$ and $\hat{\sigma}^{\mathrm{SK}}(\mathbf{x})^{2}_{\_}n$ denote the moments of that NOAA-GLM surrogate (Eq. 11), combining Sections 3.23.3, but fit to the entire campaign of runs across all 31 ensemble members and all training days/horizons/depths under study: ($\mathbf{X}_{\_}{n_{\_}M},\mathbf{Y}_{\_}{n_{\_}M}$), and $n_{\_}M\approx$ 8 million-odd runs. Providing $\mathbf{x}\in\mathcal{X}$ representing any time point(s) of interest (future, present or past), including depths and horizons of interest, yields a prediction (or forecast) capturing uncertainty in GLM dynamics driven by NOAA ensemble hindcasts. Recall from Section 2.2 that estimating a bias correction requires measuring discrepancy between computer model predictions and field data observations, forming a data set $D^{b}$ that can be used to train a GP. Our field data are a single set of observations $\mathbf{Y}^{F}_{\_}{n_{\_}F}$ in time and depth, coded in $\mathbf{X}_{\_}{n_{\_}F}$. In Section 2.2 we had $\mathbf{X}_{\_}{n_{\_}F}=\mathbf{X}_{\_}{n_{\_}M}$, so $\mathbf{Y}^{F}_{\_}{n_{\_}F}$ could easily be lined up with predictions $\boldsymbol{\mu}_{\_}{n_{\_}M}$ to measure discrepancies (6). However, now our set of GLM inputs (and outputs) is expanded [Section 3.1] to encompass horizon, which is not relevant to our field sensor measurements. Field data and surrogate are out of alignment.
@@
Recall that $\hat{\mu}_{\_}n^{\mathrm{SK}}(\mathcal{X})$ is the predictive mean of the first moments’ fit $\mathcal{GP}(D_{\_}n^{(m)})$. As in Section 2.2, $\sigma_{\_}n^{2}(\mathcal{X})$ over-estimates the variance of the sum of two GP variances because it is missing a negative covariance term. In contrast to the discussion in Section 2.2, we cannot utilize full predictive covariance matrices $\Sigma(\mathcal{X})$ because the Vecchia implementation we use only provides the diagonal, which is what we are denoting as $\sigma_{\_}n^{2}(\mathcal{X})$. The most useful settings for $\mathcal{X}$ indicate hindcasts ($\mathcal{X}$ depicting a historical set of forecasting days, horizons, and depths) and forecasts ($\mathcal{X}$ denoting the current day, a horizon of the next thirty days, and all depths of interest).
@@
To illustrate, Figure 6 shows examples of 1m-depth “slices” of hindcasts via quantities defined above starting on Feb. 19, 2021 (left), 2022 (middle) and 2023 (right). Each thin gray line in the top row is a NOAA-GLM hindcast over a 30-day horizon. Thicker lines summarize the computer model surrogate and 90% PIx (via $\hat{\mu}_{\_}n^{\mathrm{SK}}$ and $\hat{\sigma}_{\_}n^{\mathrm{SK}}$ without dividing by $n_{\_}i$), denoted as GPGLM. Darker dots show the $\bar{\mathbf{Y}}^{F}_{\_}{n_{\_}F}$ corresponding to that date/depth and each horizon. Discrepancy data $D^{b}$, derived from the $\hat{\mu}_{\_}n^{\mathrm{SK}}$ and $\bar{\mathbf{Y}}^{F}_{\_}{n_{\_}F}$, are shown in the middle panel. Observe that in all three panels/years, these fits exhibit a cold bias because the bias correction is positive. In 2021 correction from zero might not be statistically significant. In 2022 and 2023 it is substantial, highly non-linear, and error-bars do not cover zero over most of the 30-day horizon.
The bottom row of the figure shows how Eq. (12) is used to combine the quantities in the top and middle row, showing the bias fit, to obtain a bias-corrected surrogate (GPBC), for which the mean and variance are given by Eq. (12). Observe that in all three years this is a good approximation to sensor measurements. Note that while PIs for the surrogate in the top panels exhibit a trumpet-like shape over increasing horizon, those in the bottom panels no longer have this pronounced shape. Note that PIs for GPGLM use $\hat{\sigma}^{\mathrm{SK}}_{\_}n(\mathbf{x})^{2}$ without division by $n_{\_}i$ while GPBC leverages $\hat{\sigma}^{\mathrm{SK}}_{\_}n(\mathbf{x})^{2}$ with division by $n_{\_}i$ as written in Eq. (11). The total variance of GPBC is obtained by adding $\hat{\sigma}^{\mathrm{SK}}_{\_}n(\mathbf{x})^{2}$ to the predictive variance from the bias GP $(\sigma_{\_}n^{2b}(\mathcal{X}))$, shown in Eq. (12). Because $\hat{\sigma}^{\mathrm{SK}}_{\_}n(\mathbf{x})^{2}\ll\sigma_{\_}n^{2b}(\mathcal{X})$ from the (homoskedastic) bias GP, the overall predictive variance of GPBC appears homoscedastic.
@@
Figure 7 shows a similar pair of views, but now in a forecasting context corresponding to the right-most column of Figure 6 (2023). The gray lines from the GLM were used to build the surrogate (along with simulations for all other days/depths in the past), but the field data (dark dots) are not observed. Thus, bias correction is based only on patterns learned on historical data. Although the overall predictive quality in Figure 7 is lower than the bottom row of Figure 6, because the former is in-sample and the latter is out-of-sample, it is clear the GPBC performs well. Visually, the bias correction is effective (panel B) compared to the surrogate on its own (panel A). This can also be quantified as follows: GPBC achieves an RMSE of 1.24C compared to 1.86C GPGLM on its own. Note that while PI coverage is lower for GPBC for this particular example, the width of the PIs of GPBC (2.23) is less than half of the PI width of the surrogate on its own (5.14).
@@ 4.1 Iterative forecasting
Our forecasting framework consists of: an initial training stage, in which many historical NOAA-GLM simulations are used to fit a GP surrogate $\mathcal{GP}(D_{\_}{n_{\_}{M}})$ and bias correction $\mathcal{GP}(D^{b}_{\_}n)$; and an iterative stage, in which new data are downloaded and a new forecast is generated from updated surrogates. Here we are using $\mathcal{GP}$ generically, since many variations are possible. The initial training phase begins with obtaining NOAA-GLM blue hindcasts for all reference dates in the historical training period. A surrogate fit to these data provides a statistical abstraction of NOAA-GLM simulations over the historical period of time. Observed residuals from sensor data are used to estimate a bias correction.The iterative phase begins with downloading NOAA ensemble forecasts over the next thirty days. Next, these are fed as driver data to the GLM, augmenting the simulation campaign from the training phase, and updating the surrogate and bias correction accordingly. The bias-corrected surrogate can then be used to make lake temperature forecasts over that thirty-day period. The process repeats the next day after observing actual temperatures on the current day from sensors at each depth. Additional detail, along with a diagrammatic depiction [Figure B3] is provided in Section B of the Supplementary Material (Holthuijzen et al., 2025).
@@ 4.2 Incorporating annual variability into forecasts
The inputs to our surrogate modeling and bias correction framework are Julian day $t\in\{1,\dots,365\}$, depth $d\in\{0,\dots,9\}$, and horizon $h\in\{1,\dots,30\}$. A consequence of this choice, especially with regard to the time input $t$, is that our forecasts will be identical for any day $t$ in any year. We have a “day of year” model. As we gather more data, forecasts will have more conservative UQ because they will have been trained on a wider diversity of observations across year for that day, and days nearby in time. In other words, they will better capture climatological variability, but they will also be less reactive to current weather patterns and other environmental variables.
Toward obtaining more dynamic forecasts, we introduce a fourth input variable, $\phi$, whose behavior is illustrated in Figure 8. To illustrate how the “with $\phi$” model differs from its “without $\phi$” counterpart, consider the left column of Figure 8, which shows examples of predictions over a 30-day horizon for March 1, 2023 at depth 1m from the surrogate only (top panel) and after bias correction (bottom). In contrast to previous visuals, we now show all three sets of NOAA-GLM simulations available up until this time point: today’s (March 1 2023) in black and the previous two years (March 1 in 2022 and 2021) in gray. Observe that the forecasts do a poor job of tracking today’s black simulations. In the top row, forecasts from the surrogate track the corpus of simulations, whereas in the bottom row their spread is narrowed by the bias correction, but in an overly conservative fashion.
Figure 8: In-sample forecasts for March 1, 2023 models without $\phi$ (“without $\phi$”; left panels, dotted lines) and those with $\phi$ (“with” $\phi$); right panels, solid lines) at 1m depth; 90% PIS are denoted by dashed lines, Sensor measurements as dark dots. GLM forecasts for March 1, 2023 are in dark gray; previous years in light gray.
The right column of Figure 8 explores the effect of adding $\phi$ as a predictor. Observe how, in contrast with the left column without $\phi$, the GLM surrogate without a bias-correction (top) provides a good fit to NOAA-GLM forecasts (black lines). When bias-correction is added (bottom), prediction intervals narrow considerably compared to the sensor measurements (dark dots).
With $\phi$ we aim to capture the unique “state” for day $t$ and year $\lambda$ [first introduced in Section 3.1], via the following calculation
@@
i.e., average sensor-measured temperatures near the surface (depths $d$ = 0 and 1 meters) over the previous five days. Temperatures near the surface are more variable than those at deeper layers, and are therefore a good indication of weather conditions. We augment design matrices $\bar{\mathbf{X}}_{\_}n$ and $\bar{\mathbf{X}}_{\_}{n_{\_}F}^{F}$ to include $\phi(t,\lambda)$ matching up with $t$ and $\lambda$. Surrogate modeling and bias correction, as described in Sections 3.23.4, are not affected by this change excepting higher input dimension and additional lengthscale(s). In particular, responses $\bar{\mathbf{Y}}_{\_}n^{M}$ and $\mathbf{Y}^{F}_{\_}{n_{\_}F}$ remain as before.
We generally find better results with the addition of $\phi$, as demonstrated in Section 5. Consequently, we shall going forward take the $\phi$ version to be the main/default modeling variation unless we mention that modeling “without $\phi$” was entertained as a contrast. In fact, all previous visuals in the paper used the version with $\phi$, because they looked better. Eq. (13) represents only one of many ways that a year-differentiating state could be calculated. It could even be treated as a latent variable [see Section E of the Supplementary Material (Holthuijzen et al., 2025)].
@@ 5.1 Validation and benchmarking
When new sensor measurements come in at midnight we have the opportunity to assess the accuracy of yesterday’s forecasts, as predicted on all thirty previous days at varying future horizons. We evaluate (1) RMSE; (2) a proper log score (Eq. (25) of Gneiting and Raftery, 2007); (3) confidence interval width and (4) coverage. For details see Section C of the Supplementary Material (Holthuijzen et al., 2025). Smaller RMSEs and larger scores are preferred.
@@
    raw 1–30 day-ahead NOAA-GLM forecasts in the testing time period. We calculated empirical means and standard deviations over all 31 ensemble members for all combinations of reference date, depth, and horizon in the test time period.
@@ 5.2 Forecasting exercise
Panel (C) summarizes empirical coverage from 90% predictive intervals over horizon. Only the methods using surrogate-assisted NOAA-GLM forecasts can achieve nominal coverage (at 90%) at any horizon. Raw-GLM dramatically under-covers. Its surrogate (GPGLM), by contrast, is able to correctly quantify uncertainty by accessing the entire corpus of simulations across the years. Although empirical coverage drops off from nominal at higher horizons, models that incorporate GLM forecasts (GPBC and GPGLM) always outperform OGP, the climatological benchmark. Additional insight into these coverage results is provided by panel (D). Observe that the PIs of raw GLM forecasts are much narrower than those from other competitors, especially at early horizons. We also know (e.g., Figure 6) that standalone GLM simulations and forecasts can exhibit pronounced bias. Taken together, this explains raw GLM’s poor scores (panel B). Also note that PIs resulting from GPGLM are twice as wide as those of GLM for all horizons. Although GPBC and GPGLM have similar PI coverage, PIs of GPBC are much narrower than those of GPGLM. In fact, at horizon 30, PI width of GPBC (2.2) is nearly half that of GPGLM (4.2). Together, results show that bias-correction not only increases accuracy (via RMSE in panel A) but also results in narrower PIs. Note that although GPBC utilizes a homoskedastic GLM surrogate, PI interval width expands only slightly over horizon because the bias correction is homoskedastic.
Section D of the Supplementary Material (Holthuijzen et al., 2025) provides additional analysis including stratification over depth, and a decomposition of accuracy and bias over day of year.